Thanks in advanced!
in which line is a mistake first made when solving the following system of equations by elimination for s?
2r + 7s = 11
r - 9s = -7

Line 1
2r + 7s = 11
2(r - 9s = -7)
Line 2
2r + 7s = 11
2r - 18s = -14
Line 3
-11s = -3
Line 4
s = 11/3

Answer by Alan3354(69443)

in which line is a mistake first made when solving the following system of equations by elimination for s?
2r + 7s = 11
r - 9s = -7

Line 1
2r + 7s = 11
2(r - 9s = -7)
Line 2
2r + 7s = 11
2r - 18s = -14 **************
------------- Subtract
25s = 25
==============================
Line 3
-11s = -3
Line 4
s = 11/3